OpenStudy (anonymous):

Suppose y varies directly with x. If y=7 when x=2, find y when x=3 A. 6/7 B. 7/6 C. 8 D. 21/2

OpenStudy (anonymous):

\[ y=ax \text{ and } 7=a \times 2 \text{, hence } a = 3.5\] \[so y = 3.5 \times 3 = 10.5=21/2\] That is D